# Large deviations for the maximum of a reducible two-type branching Brownian motion

### v2

@@ 3.1. Upper bound
(30) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right],$
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$, and
@@
Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Furthermore, by Jensen’s inequality and Fubini’s theorem, we have
(31) $\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
(32) $\leq\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s.$
@@
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(-\varepsilon t)}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|\geq 1\right]$
@@
$\leq\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)+\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}<-\varepsilon t\right).$
Therefore,
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq 1-\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right){\rm d}s$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
@@
$\lim_{\varepsilon\rightarrow 0}\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\geq 1-\sigma^{2}\beta\theta^{2},$
@@
$\liminf_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=1-\sigma^{2}\beta\theta^{2}$
$\geq\beta(1-\theta^{2}),$
@@
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2\beta\sigma^{2}}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then for $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right]$
$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|=0\right).$
Notice that
$\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)=\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right).$
@@
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
$=\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right){\rm d}s.$
Since
$\lim_{t\rightarrow\infty}\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)=1$
@@
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right)=\beta(1-\theta^{2}),$
uniformly for $s\in[0,1]$, then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)=\beta(1-\theta^{2}).$
3.2.3. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
Notice that where $f_{1}^{\prime}(u_{0})=0$. This corresponds to the case that $\beta>1,\sigma^{2}<1$, and
We also recall that
$f(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
For $\varepsilon>0$ small enough, choose
$I_{t}=[u_{0}t,u_{0}t+1],\quad x_{t}=x_{0}t,$
for some $0<x_{0}<\sqrt{2\beta\theta^{2}\sigma^{2}}$, whose value will be determined later. Then for $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right]$
$\quad+\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right)$
$\leq\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)$
$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right),$
where
$\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0,\quad\text{and}\quad x_{0}=\frac{\sqrt{2\beta}\theta u_{0}\sigma^{3}}{1-u_{0}+u_{0}\sigma^{2}}.$
@@
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\right)$
×$\qquad\quad\times\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
We also notice that
$\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,$
since
$\sigma^{2}\geq\frac{\beta}{2\beta-1}>2-\beta,$
for $\beta>1$. Then applying (3) again, we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)$
$\quad\geq\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1.$
Since $\varepsilon>0$ is arbitrary, then applying Jensen’s inequality and Fatou’s lemma, one has
(33) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(34) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
@@
(36) $=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}$
(37) $\geq\beta(1-\theta^{2})=f_{1}(1),$
where the last inequality holds for the case that $\beta>1,\sigma^{2}<1$, and
This proves Theorem 1.1 for the above case.
@@
The arguments are exactly the same to those in Section 3. One just to replace $\sqrt{2\beta\sigma^{2}}\theta$ by $\sqrt{2}\theta$. We only give an outline here.

### v4

@@ 3.1. Upper bound
(14) $=\alpha t^{3/2}\int_{0}^{1}{\rm d}s\int_{-\infty}^{+\infty}e^{\beta st}\,\bar{u}\left((1-s)t,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-xt\right)\frac{1}{\sqrt{2\pi\sigma^{2}s}}e^{-\frac{x^{2}t}{2\sigma^{2}s}}{\rm d}x.$
By (3), for any $10<s<$ 1 and $x<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-s)$, we have, as $t\rightarrow\infty$,
$\bar{u}\left((1-s)t,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-xt\right)\leq O(1)\frac{1}{\sqrt{t}}\exp{\left\{\left(1-s-\frac{(\sqrt{2\beta\sigma^{2}\theta^{2}}-x)^{2}}{2(1-s)}\right)t\right\}}.$
Thus
$I_{2}\leq O(1)t\int_{0}^{1}{\rm d}u\int_{-\infty}^{+\infty}e^{f_{1}^{*}(u,x)t}\frac{1}{\sqrt{u}}{\rm d}x,$
where for $0<u<1,\,x\in{\mathbb{R}},$
$f^{*}_{1}(u,x)=\beta u-\frac{x^{2}}{2\sigma^{2}u}+\left(1-u-\frac{(\sqrt{2\beta\sigma^{2}\theta^{2}}-x)^{2}}{2(1-u)}\right){\bf 1}_{\{x<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u)\}},$
By studying function $x\mapsto f^{*}_{1}(u,x)$, one can see that
(15) $\max_{x\in{\mathbb{R}}}f^{*}_{1}(u,x)$ $=$ $f^{*}_{1}(u,x^{*}(u))=:f_{1}(u)$
(16) $=$ $\begin{cases}(\beta-1)u+1-\frac{\beta\sigma^{2}\theta^{2}}{1-u+\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\beta u-\frac{(\sqrt{\beta\sigma^{2}\theta^{2}}-(1-u))^{2}}{\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1,\end{cases}$
with
$x^{*}(u)=\begin{cases}\frac{\sqrt{2\beta}\theta\sigma^{3}u}{1-u+\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u),\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1,\end{cases}$
where $u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1$ is equivalent to $\frac{\sqrt{2\beta}\theta\sigma^{3}u}{1-u+\sigma^{2}u}<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u).$ Furthermore,
$f^{\prime}_{1}(u)=\begin{cases}\beta-1+\frac{\beta\sigma^{2}\theta^{2}(\sigma^{2}-1)}{(1-u+\sigma^{2}u)^{2}},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\frac{(\beta\sigma^{2}-1)u^{2}+(\sqrt{\beta\sigma^{2}\theta^{2}}-1)^{2}}{\sigma^{2}u^{2}},\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1.\end{cases}$
Also notice that for $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, we always have
$\beta\sigma^{2}\geq\begin{cases}1,\quad&\beta\leq 1,\cr\frac{\beta^{2}}{2\beta-1}\geq 1,\quad&\beta>1.\end{cases}$
Then we have two cases.
Case I. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$ for $10<u<$ 1. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
$\max_{0\leq u\leq 1,\,x\in\mathbb{R}}f^{*}_{1}(u,x)=f_{1}\left(1\right)=\beta(1-\theta^{2}).$
Hence
(17) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq\beta\left(1-\theta^{2}\right).$
Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f_{1}^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
(18) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq\beta\left(1-\theta^{2}\right).$
(19) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq 1-\theta^{2}\beta\sigma^{2}.$
Last, if $\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, then one can check that $f_{1}^{\prime}(0)\geq 0$ and $f_{1}^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$ with
(20) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\sigma\theta\cdot\sqrt{\beta}-1}{\sigma^{2}-1}.$
Therefore,
(21) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
This proves the upper bound.
Remark. Notice that
$\operatorname{Var}\left(\sigma B_{s}+B_{t}-B_{s}\right)=\sigma^{2}\cdot s+(t-s).$
Then applying Lemma 2.3 and (5), we also have
(22) $I_{2}$ $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\mathrm{d}s$
(23) $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\beta\sigma^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
(24) $=\alpha\int_{0}^{1}e^{f_{1}(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u.$
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, $\beta\leq 1$ and $\theta^{2}\leq\frac{(1-\beta)\sigma^{2}}{(\sigma^{2}-1)\beta}$, then one can check that $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$. So we will get
$\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq f^{\prime}(u_{0}),$
which is larger than $\beta(1-\theta^{2}).$ Thus we would get a rough upper bound. ∎
3.2. Lower bound for Theorem 1.1
Notice that
(25) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(26) $\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\right\}.$
First, from (3), we have
(27) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(28) $=\beta\left(1-\theta^{2}\right).$
Recall (10). Then
(29) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(30) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
(31) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right],$
where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$, and
@@
Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Then, applying Jensen’s inequality and Fubini’s theorem, gives
(32) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(33) $\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s.$
@@
$\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|\geq 1\right]$
@@
$=\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)+\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}<-\varepsilon t\right).$
This, together with (32), gives
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right){\rm d}s$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
@@
$\lim_{\varepsilon\rightarrow 0}\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\geq 1-\beta\sigma^{2}\theta^{2},$
@@
uniformly for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma again, one has
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)=1-\beta\sigma^{2}\theta^{2}\geq\beta(1-\theta^{2}),$
@@
One can make use (25) and (27) directly to conclude. We are done.∎
3.2.3. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
This corresponds to the case that $\beta>1,\sigma^{2}<1$, and
Recall $u_{0}$ from (20). We also recall that
$f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}\geq f_{1}(1)=\beta(1-\theta^{2}).$
@@
$I_{t}=[u_{0}t,u_{0}t+1],\quad x_{t}=x_{0}t,$
for some $0<x_{0}<\sqrt{2\beta\sigma^{2}\theta^{2}}$, whose value will be determined later. Then for $\varepsilon>0$ small enough and $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)\right\}\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{0}t)|\geq 1\right)+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{0}t)|=0\right),$
where
$x_{0}=\frac{\sqrt{2\beta}\theta u_{0}\sigma^{3}}{1-u_{0}+u_{0}\sigma^{2}}.$
@@
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)\right\}\right)$
×$\qquad\quad\times\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{0}t)|\geq 1\right){\rm d}s.$
Furthermore, the fact that $f_{1}^{\prime}(u_{0})=0$ gives
$1-u_{0}+u_{0}\sigma^{2}=\sqrt{\frac{1-\sigma^{2}}{\beta-1}}\sqrt{\beta}\sigma\theta.$
Then
$x_{0}=\sqrt{2}\sigma^{2}\sqrt{\frac{\beta-1}{1-\sigma^{2}}}u_{0}>\sqrt{2\beta}\sigma u_{0}$
@@
$\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}=\sqrt{2\beta}\sigma\theta\frac{1-u_{0}}{1-u_{0}+u_{0}\sigma^{2}}=\sqrt{2}\sqrt{\frac{\beta-1}{1-\sigma^{2}}}(1-u_{0})>\sqrt{2}(1-u_{0}),$
because of
$\sigma^{2}\geq\frac{\beta}{2\beta-1}>2-\beta,\quad\text{for }\beta>1.$
@@
$\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,\quad\text{for }\beta>1.$
Thus applying (3) again, we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)$
$\quad\geq\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
uniform for $s\in[0,1]$ and
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{0}t)|\geq 1\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{u_{0}t+s}>x_{0}t\right)$
$=\left(1-\frac{x_{0}^{2}}{{2\beta}\sigma^{2}u_{0}^{2}}\right)\beta u_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
uniform for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma, one has
(34) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
(35) $\geq\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}+\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
@@
(37) $=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
This proves Theorem 1.1 for the case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$. We are done.
@@
One can use arguments similar to those in Section 3. We just replace $\sqrt{2\beta\sigma^{2}}\theta$ by $\sqrt{2}\theta$. However, for the upper bound we have a simpler method.